# Robust estimation of boundary using doubly stochastic scaling of Gaussian kernel

### v2

@@ 4. Experiments
In Sections 4.1 and 4.2, we estimate points located on and near the boundary for data uniformly sampled from two geometries: (a) a circular annulus and (b) a curved truncated torus. These estimations are conducted under three noise scenarios: no noise, homoskedastic noise, and heteroskedastic noise. Additionally, in Section 4.3, we identify images in the MNIST dataset that lie close to and far from the boundary. We evaluate the performance of the following methods for boundary estimation:
@@
    The doubly stochastic kernel with local PCA as in Algorithm 1, and a variant of it which, while calculating $\bm{\nu}_{i}$ in Eq. 35, directly uses the difference in the observations $\widetilde{\mathbf{x}}_{j}-\widetilde{\mathbf{x}}_{i}$ as in Eq. 31, instead of the local PCA projections. These methods are referred to as DS+LPCA and DS, respectively.
@@
    The method proposed in [6] that estimates the boundary using an approach similar to Gaussian, but replaces the Gaussian kernel with a binary cutoff kernel. We refer to this method as Binary. The implementation of this method is available in [5].
@@
(40) $\mathcal{B}_{p}=\{\mathbf{x}_{k}:b_{\mathbf{x}_{k}}\leq p\text{-th percentile of }(b_{\mathbf{x}_{i}})_{1}^{n}\}$
and the estimated $p$-th percentile boundary as
(41) ^$\widehat{\mathcal{B}}_{p}=\begin{cases}\{\mathbf{x}_{k}:\widehat{b}_{\mathbf{x}_{k}}\leq p\text{-th percentile of }(\widehat{b}_{\mathbf{x}_{i}})_{1}^{n}\}\text{ with {{Gaussian}}, {{Gaussian+LPCA}} \& {{Binary}}}\\ \{\mathbf{x}_{k}:\left\|\bm{\nu}_{k}\right\|\geq(100-p)\text{th percentile of }(\left\|\bm{\nu}_{i}\right\|)_{1}^{n}\}\text{ with {{DS}} \& {{DS+LPCA}}}.\end{cases}$
To measure the discrepancy between the estimated and true boundaries, we use the Jaccard index,
@@
4.1. Data descriptions
For our first set of datasets, we generate $n=7076$ uniformly distributed points in $\mathbb{R}^{2}$, sampled as $\{(r_{i}\cos(\theta_{i}),r_{i}\sin(\theta_{i}))\}_{1}^{n}$ from a circular annulus with an outer radius of $1$ and an inner radius of $0.3$, i.e. $r_{i}\in[0.3,1]$ and $\theta_{i}\in(-\pi,\pi]$ (see Figure 2). We then project these points into $m=2000$ dimensions using an orthogonal transformation. This forms our noiseless dataset. Let $\mathcal{S}$ be the $2$-dimensional subspace containing the annulus. We construct two noisy variants of the data (a) by adding homoskedastic noise in a ball of radius $0.2$ at each point, ensuring that the noise is orthogonal to $\mathcal{S}$ (this preserves the ground truth distances of the points from the boundary for evaluation) and (b) by adding heteroskedastic noise in a ball of radius $20.01+0.2(1+\cos(2\theta_{i}))/$ 2 at each point, also orthogonal to $\mathcal{S}$. The noise distribution is illustrated in Figure 2.
@@ 4.1. Data descriptions
on a curved torus in $\mathbb{R}^{3}$ where $R=0.25$, $r=1/4\pi^{2}R$, $\theta_{i}\in[0,2\pi)$ and $\phi_{i}\in[0,2\pi)$. The torus is then truncated by removing points with $x$-coordinate greater than $0.18$, as shown in Figure 3. These points are projected into $m=2000$ dimensions using an orthogonal transformation, forming our noiseless dataset. Let $\mathcal{S}$ represent the $3$-dimensional subspace containing the truncated torus. We generate two noisy variants of the data (a) by adding homoskedastic noise in a ball of radius $0.075$ at each point while ensuring that the noise is orthogonal to $\mathcal{S}$ and (b) by adding heteroskedastic noise in a ball of radius $20.01+0.075(1+\cos(2\theta_{i}))/$ 2 at each point, also orthogonal to $\mathcal{S}$. The noise distribution is illustrated in Figure 3.
@@
4.2. Interpreting results
For both datasets, we construct the standard Gaussian kernel and the doubly stochastic kernel [17] using a bandwidth of $h=0.1$. Using these kernels, $d=2$ and $k_{\text{nn}}=256$ neighbors for local PCA computations, we estimate the distances to the boundary using DS+LPCA, DS, Gaussian+LPCA and Gaussian. We use the same number of nearest neighbors while estimating distances to boundary using Binary. Subsequently we compute the $p$-th percentile boundary ^$\widehat{\mathcal{B}}_{p}$ (Eq. 41) for each method. To compare the performance, we calculate the Jaccard index $\mathcal{J}_{p}$ (Eq. 42) against the ground truth boundary $\mathcal{B}_{p}$ (Eq. 40) for various values of $p$. The results are presented in Figures 4, 5, and 6, and we summarize key observations from the figures below.
@@ 4.2. Interpreting results
      The curvature of the torus makes local PCA an important factor. We observe that DS+LPCA and Gaussian+LPCA achieve comparable performance and outperform DS and Gaussian, indicating that local PCA enhances the accuracy of boundary estimation. At the same time, the performance of Binary degrades significantly, suggesting the method is sensitive to the curvature. This is consistent with the observation in [6] that the method Binary is designed for Euclidean domains, not general smooth manifolds.
@@
    Under homoskedastic noise: For both the circular annulus and the truncated torus, the Jaccard index for DS+LPCA, DS, and Gaussian+LPCA remains robust against noise, whereas the performance of Gaussian and Binary deteriorates significantly. The degradation of Gaussian can be attributed to the behavior of the BDE observed in Figure 6 where the BDE from Gaussian is completely corrupted by noise. Specifically, the BDE is relatively large in the interior and smaller at the boundary, contrary to the expected trend as illustrated in Figure 1. In fact, the norm of the BDE is so large that the estimated $\widehat{b}_{\mathbf{x}_{i}}$ collapse to zero (due to Newton’s method converging to negative values that are clipped to zero), resulting in degenerate boundary in Figure 4.
    Interestingly, while the norm of the BDE for DS increases in the interior compared to the noiseless case, the boundary values remain relatively higher. This ensures that the estimates $\left\|\bm{\nu}_{i}\right\|$ at the boundary are higher than those in the interior, enabling more accurate estimates of boundary points. Lastly, as shown in Figure 6, DS+LPCA and Gaussian+LPCA exhibit similar BDE values, leading to comparable boundary estimates as shown in Figure 4, consistent with the noiseless case. This again highlights that incorporating local PCA significantly enhances the accuracy of boundary estimation, as evidenced by the substantial increase in the Jaccard index from Gaussian to Gaussian+LPCA.
@@ 4.3. MNIST
In this experiment, we start by randomly selecting $5000$ images of size $28\times 28$ for each of the $10$ digits in the MNIST dataset, resulting in $10$ datasets, each containing $n=5000$ samples in $m=784$ dimensions. For each dataset, the median distance to the $128$th nearest neighbor of each image is used as the bandwidth $h$, and the local PCA is computed using $k_{\text{nn}}=512$ and $d=2$. Using DS, DS+LPCA, and Gaussian+LPCA, we compute the $10$th percentile boundary, ^$\widehat{\mathcal{B}}_{10}$, as defined in Eq. 41. These results are visualized on a two-dimensional UMAP [21] embedding of the entire dataset in Figure 7. It is important to note that the boundary estimation is performed on the raw datasets, not the UMAP embeddings.
@@
Consistent with our simulations on synthetic datasets, the boundary estimated by DS and Binary are highly noisy, with a Jaccard index of $0.257$ and $0.083$ respectively, when compared to the boundary estimated by DS+LPCA. Note that there is no ground-truth boundary in this context, and unlike the scenarios in previous sections, the Jaccard index here serves to quantify the qualitative differences between the methods. From Figure 7, it is evident that DS and Binary misclassify numerous points as being close to the boundary. In contrast, the boundaries computed using DS+LPCA and Gaussian+LPCA have a significant overlap, with a Jaccard index of $0.847$. However, there are subtle differences. As shown in Figure 7, points that lie on the boundary estimated by DS+LPCA but not on the one by Gaussian+LPCA tend to be more dispersed along the boundary, while Gaussian+LPCA identifies points that are more concentrated in tightly packed regions resembling corners in the embedding, as is evident for digits $3$, $5$, $7$, and $9$. This suggests that Gaussian+LPCA provides a skewed boundary estimate than DS+LPCA, resulting in boundary points localized within specific regions.
@@
Finally, to examine the differences between images near and far from the boundary, we randomly sampled $15$ images from ^$\widehat{\mathcal{B}}_{5}$ and outside ^$\widehat{\mathcal{B}}_{95}$, as estimated using DS+LPCA. The former set represents images near the boundary, while the latter represents images in the interior. These images are visualized in Figure 8. Clearly, images within the interior exhibit less variation, whereas those near the boundary display greater variation. The contrast is evident in the average of the $15$ images per digit, displayed in the last columns of Figure 8. This highlights that the boundary points capture more diverse and nuanced features of the dataset, emphasizing the importance of accurately identifying them. Furthermore, this insight opens new avenues for exploring the computational benefits and generalization performance of machine learning models trained solely on boundary points rather than on the entire dataset.
@@ 5. Conclusion
In this work, we have developed a robust framework for detecting boundary points for data sampled from a manifold with boundary and in the presence of noise. To accomplish this, we extended the recent results in doubly stochastic scaling of the Gaussian kernel [20, 18, 17, 8] on closed manifolds to the setting of manifolds with boundaries. Specifically, we obtained a novel characterization of the scaling factors that transform a Gaussian kernel into a doubly stochastic kernel on manifolds with boundary. We further developed a BDE leveraging the doubly stochastic kernel along with local PCA, and utilized our characterization of scaling factors to prove its convergence.
Our experiments demonstrated that our proposed BDE significantly improves the detection of boundary points compared to the standard Gaussian and binary kernel-based methods [1, 6], particularly in noisy conditions (Figures 4, 5, and 6). Our simulations underline the importance of incorporating both doubly stochastic scaling and local PCA for robust performance. Additionally, we observed that boundary points capture a wider range of features within the data (Figure 8). In future work, we aim to assess the trade-off between training models exclusively on boundary points versus the full dataset. This includes exploring the impact on model performance, computational efficiency, and generalization capabilities.
6. Proofs

### v3

@@ 5. Experiments
In Section 5.1, we estimate points located on and near the boundary for data uniformly sampled from two geometries: (a) a circular annulus and (b) a curved truncated torus. These estimations are conducted under three noise scenarios: no noise, homoskedastic noise, and heteroskedastic noise. In Section 5.2, we demonstrate how our boundary correction term improves density estimation on the circular annulus and truncated torus under heteroskedastic noise. Finally, in Section 5.3, we apply our method to high-dimensional MNIST dataset to identify images that lie close to and far from the boundary. In the following experiments, we evaluate the performance of the following methods:
@@
    Our approach using the doubly stochastic kernel with local PCA as in Algorithm 1, and a variant of it which, while calculating $\bm{\nu}_{i}$ in Eq. 37, directly uses the difference in the observations $\widetilde{\mathbf{x}}_{j}-\widetilde{\mathbf{x}}_{i}$ as in Eq. 33, instead of the local PCA projections. These methods are referred to as DS+LPCA and DS, respectively.
@@
    The method proposed in [calder2022boundary] that estimates the boundary using an approach similar to Gaussian, but replaces the Gaussian kernel with a binary cutoff kernel. We refer to this method as Binary. The implementation of this method is available in [graphlearning].
@@
(43) $\mathcal{B}_{p}=\{k:b_{\mathbf{x}_{k}}\leq p\text{-th percentile of }(b_{\mathbf{x}_{i}})_{1}^{n}\}.$
For the methods Gaussian, Gaussian+LPCA and Binary, the estimated $p$-th percentile boundary is set to
(44) $\widehat{\mathcal{B}}_{p}=\{k:\widehat{b}_{\mathbf{x}_{k}}\leq p\text{-th percentile of }(\widehat{b}_{\mathbf{x}_{i}})_{1}^{n}\},$
and for DS and DS+LPCA, it is set to
(45) $\widehat{\mathcal{B}}_{p}=\{k:\left\|\bm{\nu}_{k}\right\|\geq(100-p)\text{th percentile of }(\left\|\bm{\nu}_{i}\right\|)_{1}^{n}\}.$
To quantify the discrepancy between the estimated and true boundaries, we use the Jaccard index,
@@
5.1. Robust boundary estimation
5.1.1. Data descriptions
For our first set of datasets, we generate $n=7076$ uniformly distributed points in $\mathbb{R}^{2}$, sampled as $\{(r_{i}\cos(\theta_{i}),r_{i}\sin(\theta_{i}))\}_{1}^{n}$ from a circular annulus with an outer radius of $1$ and an inner radius of $0.3$, i.e. $r_{i}\in[0.3,1]$ and $\theta_{i}\in(-\pi,\pi]$ (see Figure 2). We then project these points into $m=2000$ dimensions using an orthogonal transformation. This forms our noiseless dataset. Let $\mathcal{S}$ be the $2$-dimensional subspace containing the annulus. We construct two noisy variants of the data (a) by adding homoskedastic noise in a ball of radius $0.2$ at each point, ensuring that the noise is orthogonal to $\mathcal{S}$ (this preserves the ground truth distances of the points from the boundary for evaluation) and (b) by adding heteroskedastic noise in a ball of radius $0.01+0.2(1+\cos(2\theta_{i}))/2$ at each point, also orthogonal to $\mathcal{S}$. The noise distribution is illustrated in Figure 2.
@@ 5.1.1. Data descriptions
on a curved torus in $\mathbb{R}^{3}$ where $R=0.25$, $r=1/4\pi^{2}R$, $\theta_{i}\in[0,2\pi)$ and $\phi_{i}\in[0,2\pi)$. The torus is then truncated by removing points with $x$-coordinate greater than $0.18$, as shown in Figure 3. These points are projected into $m=2000$ dimensions using an orthogonal transformation, forming our noiseless dataset. Let $\mathcal{S}$ represent the $3$-dimensional subspace containing the truncated torus. We generate two noisy variants of the data (a) by adding homoskedastic noise in a ball of radius $0.075$ at each point while ensuring that the noise is orthogonal to $\mathcal{S}$ and (b) by adding heteroskedastic noise in a ball of radius $0.01+0.075(1+\cos(2\theta_{i}))/2$ at each point, also orthogonal to $\mathcal{S}$. The noise distribution is illustrated in Figure 3.
@@
5.1.2. Interpreting results
For both datasets, we construct the standard Gaussian kernel and the doubly stochastic kernel [landa2023robust] using a bandwidth of $h=0.1$. Using these kernels, $d=2$ and $k_{\text{nn}}=256$ neighbors for local PCA computations, we estimate the distances to the boundary using DS+LPCA, DS, Gaussian+LPCA and Gaussian. We use the same number of nearest neighbors while estimating distances to boundary using Binary. Subsequently we compute the $p$-th percentile boundary $\widehat{\mathcal{B}}_{p}$ for each method (Eq. 44-45). To compare the performance, we calculate the Jaccard index $\mathcal{J}_{p}$ (Eq. 46) against the ground truth boundary $\mathcal{B}_{p}$ (Eq. 43) for various values of $p$. The results are presented in Figures 4, 5, and 6, and we summarize key observations from the figures below.
@@ 5.1.2. Interpreting results
      The curvature of the torus makes local PCA an important factor. We observe that DS+LPCA and Gaussian+LPCA achieve comparable performance and outperform DS and Gaussian, indicating that local PCA enhances the accuracy of boundary estimation. At the same time, the performance of Binary degrades significantly, suggesting the method is sensitive to the curvature. This is consistent with the observation in [calder2022boundary] that the method Binary is designed for Euclidean domains, not general smooth manifolds.
    Circular Annulus
    Figure 4. (Top) The Jaccard index (Eq. 46) is shown between the ground truth $p$-th percentile boundary $\mathcal{B}_{p}$ and the estimated $p$-th percentile boundary $\widehat{\mathcal{B}}_{p}$ for $p\in[2,14]$. (Bottom) Visualizations of the boundaries $\mathcal{B}_{p}$ and $\widehat{\mathcal{B}}_{p}$ (in orange) are provided for competing methods at $p\in\{2,5,8,11\}$.
    Truncated torus
    Figure 5. (Top) The Jaccard index (Eq. 46) is shown between the ground truth $p$-th percentile boundary $\mathcal{B}_{p}$ and the estimated $p$-th percentile boundary $\widehat{\mathcal{B}}_{p}$ for $p\in[2,14]$. (Bottom) Visualizations of the boundaries $\mathcal{B}_{p}$ and $\widehat{\mathcal{B}}_{p}$ (in orange) are provided for competing methods at $p\in\{2,5,8,11\}$. The visualizations are presented on the two-dimensional domain of the truncated torus.
@@
    Under homoskedastic noise: For both the circular annulus and the truncated torus, the Jaccard index for DS+LPCA, DS, and Gaussian+LPCA remains robust against noise, whereas the performance of Gaussian and Binary deteriorates significantly. The degradation of Gaussian can be attributed to the behavior of the BDE observed in Figure 6 where the BDE from Gaussian is completely corrupted by noise. Specifically, the BDE is relatively large in the interior and smaller at the boundary, contrary to the expected trend as illustrated in Figure 1. In fact, the norm of the BDE is so large that the estimated $\widehat{b}_{\mathbf{x}_{i}}$ collapse to zero (due to Newton’s method converging to negative values that are clipped to zero), resulting in degenerate boundary in Figure 4. Interestingly, while the norm of the BDE for DS increases in the interior compared to the noiseless case, the boundary values remain relatively higher. This ensures that the estimates $\left\|\bm{\nu}_{i}\right\|$ at the boundary are higher than those in the interior, enabling more accurate estimates of boundary points. Lastly, as shown in Figure 6, DS+LPCA and Gaussian+LPCA exhibit similar BDE values, leading to comparable boundary estimates as shown in Figure 4, consistent with the noiseless case. This again highlights that incorporating local PCA significantly enhances the accuracy of boundary estimation, as evidenced by the substantial increase in the Jaccard index from Gaussian to Gaussian+LPCA.
    Circular Annulus Truncated torus
    Refer to caption Refer to caption
    Figure 6. (Top) The theoretical norm of the BDE (Eq. 35 and Eq. 38) at each point on the data manifold are visualized. These are computed using the true distances to the boundary and uniform density. (Bottom) The empirical norm of the BDE due to each method (Eq. 33 and Eq. 37) are visualized on the data manifold. For each case, the values are normalized by the maximum. For the truncated torus, the visualizations are presented onto its two-dimensional domain.
@@
5.2. Boundary-corrected density estimation
5.2.1. Datasets
Once again, we sampled points $\{(r_{i}\cos(\theta_{i}),r_{i}\sin(\theta_{i}))\}_{1}^{n}$ from a circular annulus with $r_{i}\in[0.3,1]$ and $\theta_{i}\in(0,2\pi]$. This time, however, we sample $n=7076$ points with non-uniform density where $\theta_{i}$ is sampled from $\text{Normal}(0,0.25\pi^{2})$ modulo $2\pi$ and $r_{i}\sim\text{Uniform}(0.3,1)$. We use the heteroskedastic noise conditions $0.01+0.2(1+\cos(2\theta_{i}))/2$ from Section 5.1.1. For the second dataset, we sample $n=11,175$ points from a truncated curved torus with heteroskedastic noise in a ball of radius $0.01+0.075(1+\cos(2\theta_{i}))/2$, this time with a non-uniform density where $\theta_{i}\sim\text{Truncated-Normal}((-\pi,\pi),0,0.75\pi)$ and $\phi_{i}\sim\text{Uniform}(0,2\pi)$.
5.2.2. Results
For the annulus, we set the intrinsic dimension to $d=2$, the bandwidth to $h=0.1$, and the number of nearest neighbors to $k_{\mathrm{nn}}=1024$ for both the doubly stochastic and standard Gaussian kernels. The same kernels are used for both boundary estimation and kernel density estimation. For the KDE defined in Eq. (41), we fix the parameters to $s=8.0$ and $u=0$; an ablation study over these parameters is provided in Figure 9. As shown in Figure 7, the KDE without boundary correction systematically underestimates the density near the boundary due to kernel mass extending beyond the support, as discussed in Section 4.3. The boundary correction term in Eq. (41) effectively resolves this bias. In contrast, the standard Gaussian kernel–based KDE with boundary correction [berry2017density] remains corrupted by heteroskedastic noise.
Figure 7. Comparison of kernel density estimators on a circular annulus with heteroskedastic noise compared to the ground truth density $q$ (left). Each estimate is normalized by its respective maximum value.
Figure 8. Comparison of kernel density estimators on truncated torus with heteroskedastic noise compared to the ground truth density $q$ (left). Each estimate is normalized by its respective maximum value.
Figure 9. The two-parameter family of kernel density estimators defined in Eq. (41), evaluated for various values of $s$ and $u$ on a circular annulus and a torus with heteroskedastic noise, in the setting where the true distances to the boundary are known a priori. As before, each estimate is normalized by its respective maximum value.
For the truncated torus, we again set $d=2$ and $h=0.1$. Boundary estimation is performed using $k_{\mathrm{nn}}=256$, after which the KDE is computed with $k_{\mathrm{nn}}=1024$ once the distances to the boundary have been estimated. As in the annulus case, we fix $s=8$ and $u=0$ in Eq. (41), with a corresponding ablation study reported in Figure 9. The results in Figure 8 again demonstrate that omitting boundary correction leads to an underestimation of the density near the boundary. In contrast to the annulus, the proposed boundary-corrected KDE in Eq. (41) exhibits a mild overestimation near the boundary; nevertheless, it consistently outperforms the standard Gaussian kernel–based KDE with boundary correction.
@@ 5.3. MNIST
In this experiment, we randomly select $5000$ images of size $28\times 28$ for each of the $10$ digits in the MNIST dataset, resulting in $10$ datasets, each containing $n=5000$ samples in $m=784$ dimensions. For each dataset, the median distance to the $128$th nearest neighbor of each image is used as the bandwidth $h$, and the local PCA is computed using $k_{\text{nn}}=512$ and $d=2$. Using DS, DS+LPCA, Gaussian+LPCA and Binary, we compute the $10$th percentile boundary, $\widehat{\mathcal{B}}_{10}$, as defined in Eq. 44-45. These results are visualized on a two-dimensional UMAP [umap] embedding of the entire dataset in Figure 10. It is important to note that the boundary estimation is performed on the raw datasets, not the UMAP embeddings. Since we expect the boundary points to be localized and perhaps supported on curve segments, their visualization via UMAP can provide a qualitative assessment of the various boundary detection algorithms. It is also worth noting that certain points which seem to lie on the boundary of the clusters produced within the UMAP embedding in Figure 10 are not detected as boundary points by either DS+LPCA or Gaussian+LPCA.
@@
Because there is no ground-truth boundary, we cannot evaluate accuracy directly as we did in previous sections. Instead, we are going use the Jaccard index to quantify the relative, qualitative agreement between the different methods. Treating DS+LPCA as a reference, we find that the boundaries estimated by DS and Binary diverge significantly, yielding Jaccard indices of $0.257$ and $0.083$, respectively. It is visually evident from Figure 10 that boundary points detected by DS and Binary are highly scattered, inconsistent without our expectation of localized boundaries. In contrast, the boundaries computed using DS+LPCA and Gaussian+LPCA are concentrated along curve segments, and have a significant overlap with a Jaccard index of $0.847$. However, there are subtle differences. As shown in Figure 10, points that lie on the boundary estimated by DS+LPCA but not on the one by Gaussian+LPCA tend to be more dispersed along the boundary, while Gaussian+LPCA identifies points that are more concentrated in tightly packed regions resembling corners in the embedding, as is evident for digits $3$, 55, 77, and $9$. This suggests that boundary estimate by Gaussian+LPCA is more skewed toward high-curvature, corner-like regions, in comparison to DS+LPCA.
@@
Finally, to examine the differences between images near and far from the boundary, we randomly sampled $15$ images from $\widehat{\mathcal{B}}_{5}$ and outside $\widehat{\mathcal{B}}_{95}$, as estimated using our DS+LPCA. The former set represents images near the boundary, while the latter represents images in the interior. These images are visualized in Figure 11. Clearly, images within the interior exhibit less variation, whereas those near the boundary display greater variation. The contrast is evident in the average of the $15$ images per digit, displayed in the last columns of Figure 11. This highlights that the boundary points capture more diverse and nuanced features of the dataset, emphasizing the importance of accurately identifying them. Furthermore, this insight opens new avenues for exploring the computational benefits and generalization performance of machine learning models trained solely on boundary points rather than on the entire dataset.
Figure 10. The $10$-th percentile boundary $\widehat{\mathcal{B}}_{10}$ (as defined in Eq. 44-45) is estimated from $5000$ raw $28\times 28$ images of each MNIST digit using DS+LPCA, Gaussian + LPCA, DS and Binary. The visualizations are presented on a two-dimensional UMAP[umap] embedding of the data, where points in $\widehat{\mathcal{B}}_{10}$ are highlighted in black. These plots show points on the estimated boundaries due to each method, the ones that lie on the estimated boundaries by two of the methods, and by one method but not the other.
Figure 11. A random selection of 15 images per digit is shown, including those from the interior of the MNIST dataset (outside $\widehat{\mathcal{B}}_{95}$) and near the boundary (inside $\widehat{\mathcal{B}}_{5}$), as estimated using DS + LPCA. The final column displays the average of the 15 images.
@@ 6. Conclusion
In this work, we developed a robust framework for detecting boundary points in data sampled from noisy manifolds with boundary. To this end, we extended recent results on the doubly stochastic scaling of Gaussian kernels on closed manifolds [marshall2019manifold, landa2021doubly, landa2023robust, cheng2024bi] to the setting of manifolds with boundary. In particular, we derived a characterization of the scaling factors that transform a Gaussian kernel into a doubly stochastic kernel in the setting of a manifold with boundary. Building on this characterization, we proposed a BDE that combines the doubly stochastic kernel with local PCA, and utilized our characterization of the scaling factors to establish its convergence.
Numerical experiments demonstrate that the proposed BDE substantially outperforms standard Gaussian and binary kernel–based methods [berry2017density, calder2022boundary] in detecting boundary points, especially in high-noise regimes (Figures 4, 5, and 6). Moreover, we observe that the detected boundary points capture a broader range of salient features within the data (Figure 11).
We further introduced a two-parameter family of kernel density estimators that extends the KDE proposed in [landa2023robust] for closed manifolds to manifolds with boundary by incorporating a boundary correction term. Our experiments (Figures 7 and 8) show that the proposed estimators yield improved density estimates compared to standard Gaussian KDEs with boundary correction.
Overall, this work highlights the importance of incorporating both doubly stochastic scaling and local PCA to achieve robust performance. As future work, we plan to investigate the trade-offs associated with training models using only boundary points versus the full dataset, with focus on on predictive performance, computational efficiency, and generalization.
Appendix A Proofs